a company had to downsize and fire 20 employees. Of these 20 employees, 10 were women. However, only 33% of the company’s entire workforce were women. The women who were fired filed a sexual discrimination lawsuit since 50% of the fired employees were women. Determine the probability that 50% or more of the fired employees are women. Then determine if the women’s claim of sexual discrimination is legitimate
Answer #1

We use a hypothesis test to decide whether the company is firing women with a higher probability that their proportion in the workforce.

Let be the indicator variable which takes value 1 if the person being fired is a woman. This is a Bernoulli variable with parameter p (where p is the probability of the fired person being woman)

Now, let be the number of women in a sample of 20 fired people . Each person is chosen to be fired independently of the rest. Then,

Use this to find (I've computed this value using a Python code and the screenshot is attached below the answer.)

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

het Ho: There is no gender discrimination p= 0-33 Hi : There is gender dicrimination P20-33 be the test where p is the probability of choosing a woman to be fired. This is a non-parametric test. we are interested in probability that knod at least 10 women are chosen to be fired out of 20 people who are to be fired assuming to (p=0.33) Let X be the number of women out of the 20. people. Then X~Binomial (2o, p) Under Ho, p=0.33. Pro (X7,10) = P6X = i) = 360) (0-3)(0.67 I=ID = 0.0065 > 0.05 (computed wing ) I Python code At 95-7. significance level, there is not enough evidence to reject Ho (ie there is not enough evidence of gender discrimination)

In [9]: from math import factorial as fac def calc_prob (n,x,pl,p2): return pl*** * P2**(n-x)*fac(n)/(fac(x)*fac (n-x)) In [13]: prob = 0 for i in range (10,21): print("Probability that "+str(i)+" of the 20 people are women = "+str(calc_prob (20,1, 0.33,0.67))) prob = prob + calc_prob (20,1,0.33, 0.67) print("\nProbability that atleast 10 of the 20 people are women = "+str(prob)) Probability that 10 of the 20 people are women = 0.051580550977615776 Probability that 11 of the 20 people are women = 0.023095769094454828 Probability that 12 of the 20 people are women = 0.008531646046085925 Probability that 13 of the 20 people are women = 0.0025859409370455615 Probability that 14 of the 20 people are women = 0.0006368362009142056 Probability that 15 of the 20 people are women = 0.00012546623659802255 Probability that 16 of the 20 people are women = 1.9311500968912057e-05 Probability that 17 of the 20 people are women = 2.2380317189608357e-06 Probability that 18 of the 20 people are women = 1.837190217057402e-07 Probability that 19 of the 20 people are women = 9.525102460784646e-09 Probability that 20 of the 20 people are women = 2.3457341881036816e-10 Probability that atleast 10 of the 20 people are women = 0.08657795250409978